### v1

# On additive error approximations to ##\##BQP

Abstract
Counting complexity characterizes the difficulty of computing functions related to the number of valid certificates to efficiently verifiable decision problems. Here we study additive approximations to a quantum generalization of counting classes known as #BQP. First, we show that there exist efficient quantum algorithms that achieve additive approximations to #BQP problems to an error exponential in the number of witness qubits in the corresponding verifier circuit, and demonstrate that the level of approximation attained is, in a sense, optimal. We next give evidence that such approximations can not be efficiently achieved classically by showing that the ability to return such approximations is BQP-hard. We next look at the relationship between such additive approximations to #BQP and the complexity class DQC1, showing that a restricted class of #BQP problems are DQC1-complete.
1 Introduction
Given a decision problem with polynomial-time verifiable solutions, a natural question to ask is how many solutions, i.e., accepting witnesses, exist. When the verifier is classical, this question essentially defines the counting complexity class #P [26]—the class of functions that can be expressed as the number of accepting witnesses to NP problems. These counting problems appear in combinatorics, graph theory, and statistical physics but are known to be extremely hard to solve. In fact, a result of Toda [25] implies that a polynomial number of queries to a #P oracle can solve any problem in the entire polynomial hierarchy.
The class #P precisely characterizes the complexity of counting problems such as computing the permanent of binary matrices, counting the number of perfect matchings in a graph [26], and computing partition functions of certain classical spin systems [17, 6]. On the other hand, a number of important problems in, e.g., algebraic combinatorics [7] and quantum computing [5] are polynomial-time reducible to #P but are not #P problems. The closure of #P under subtraction defines the class GapP [12] which is known to contain problems such as computing multiplicities arising in representation theory of the symmetric group [7].
The counting class #BQP was defined to formalize the notion of counting the number of quantum witnesses to promise problems that allow efficient verification on quantum computers [5, 21]. This class contains quantum-inspired counting problems such as computing the density of states of quantum Hamiltonians [5] and computing representation theoretic multiplicities such as Kronecker coefficients [16, 3]. Though #P, GapP, and #BQP are equivalent under polynomial-time reductions [5, 11], the subtle difference amongst them manifests directly in the hardness of approximation, specifically approximating the respective quantities to within a multiplicative error. Any #P function can be multiplicatively approximated in the third level of the polynomial hierarchy [24], a class of problems believed to be easier than #${\sf\#P}$, but approximating GapP functions can be #P-hard in general [14]. However, it is unknown if multiplicatively approximating #BQP relations is any easier even though this has been used to define a new class QXC of quantum approximate counting problems [3, 4].
It is thus interesting to ask if other notions of approximation also lead to problems of varying complexity. The study of additive approximations to counting problems was initiated in Ref. [1], motivated by the result that certain additive approximations to GapP problems suffice to solve BQP decision problems [12]. There, it was shown that all #P and GapP functions have efficient classical additive approximation algorithms up to a normalization factor exponential in the size of the descriptions of the solutions, with no difference in the complexity of obtaining the additive approximations for #P and GapP. In quantum computation, additive error approximations to some #BQP and QXC problems have been studied in the context of DQC1, the class of decision problems that can be solved in a model of quantum computation where the system is initialized with only one pure qubit [18, 20]. Specifically, decision problems based on certain additive error approximations to the density of states and quantum partition functions have been shown to be complete for DQC1 [2, 9, 8].
In this work, we study the complexity of obtaining additive error approximations to #BQP problems and its connection to the complexity classes BQP and DQC1. We prove that there exists an efficient quantum algorithm which produces additive approximations to all #BQP problems up to a normalization exponential in the the number of qubits of the quantum witness. We then give evidence that the level of approximation achieved quantumly is not achievable by any polynomial-time classical algorithm by showing that it is BQP-hard. Next, we show that this level of approximation is likely optimal as the problem of computing more accurate approximations becomes #BQP-hard. We demonstrate, however, that a weaker additive approximation is achievable via a polynomial-time classical algorithm. Next we formalize results that elucidate the connection between DQC1 and #BQP. We show that DQC1 is equivalent to computing additive approximations to a subset of #BQP problems, namely those with verifier circuits containing at most a logarithmic number of ancilla qubits.
2 Background and Summary of Results
2.1 Classical Counting Complexity
Classical counting complexity, including classes such as #P and GapP, aims to characterize the difficulty of computing functions related to the number of accepting witnesses to NP problems. A decision problem is in NP if given an input $x$ that is a yes instance, there exists an efficiently generated polynomial-time verification procedure and at least one witness $y$ such that the procedure will accept on the input $(x,y)$, while for no instances, no such witness exists. The class #P is the set of functions that can be defined by counting the number of witnesses to an NP problem.
We can conveniently define both NP and #P via relations $R\subset\{0,1\}^{*}\times\{0,1\}^{*}$ between the set of inputs and the set of possible solutions such that there exists a Turing machine that decides the proposition $R(x,y)$ in polynomial-time. Introducing the set of accepting witnesses (of length bounded by a polynomial $p$ of the input) for input $x$,
$R(x)=\{y:R(x,y),|y|\leq p(|x|)\}\,,$ (1)
solving the NP decision problem for input $x$ reduces to deciding whether $|R(x)|>0$ while solving the #P problem for input $x$ is to return $|R(x)|$. The class GapP is then defined as the closure of #P under subtraction. Namely, a function is in GapP if and only if it can be expressed as the difference of two #P functions.
2.2 Acceptance Operators, QMA, and DQC1
A quantum generalization of NP is the class QMA, which roughly corresponds to NP except the witnesses are quantum states and the verification procedure is a quantum circuit. We will define QMA, and later the quantum analogue of #P, known as #BQP, via properties of acceptance operators $\mathcal{V}$ [4, 5] which generalize the relations $R$ in the previous section. We define them via a quantum circuit $V$ that acts on a tensor product Hilbert space $\mathcal{H}_{I}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{W}$ where $\mathcal{H}_{W}$ is the subsystem corresponding to the quantum witnesses, $\mathcal{H}_{A}$ corresponds to the subsystem of the ancillas needed for the verification procedure and $\mathcal{H}_{I}$ is the subsystem corresponding to the (classical) input. Given the circuit $V$ requiring $a$ ancilla qubits and a classical input bit string $x$, we define the acceptance operator $\mathcal{V}_{x}$ via
$\mathcal{V}_{x}=(\bra{x}\otimes\bra{0^{a}}\otimes\mathbbm{I}_{W})V^{\dagger}\Pi_{\text{out}}V(\ket{x}\otimes\ket{0^{a}}\otimes\mathbbm{I}_{W})\,,$ (2)
where $\Pi_{\text{out}}$ is a projector onto the |$\outerproduct{1}{1}$ | state of the first qubit. With this we define the complexity class QMA.
Definition 1 (QMA).
Let $A=(A_{\text{YES}},A_{\text{NO}})$ be a promise problem and let $c,s:\mathbb{N}\to[0,1]$ be functions satisfying $c(n)-s(n)=\Omega(\text{poly}(n)^{-1})$. Then $A\in{\sf QMA}$ if and only if there exists a polynomial-time uniform family of quantum circuits $\{{V^{(n)}}:n\in\mathbb{N}\}$ acting on $w(n)=O(\text{poly}(n))$ witness qubits, $a(n)=O(\text{poly}(n))$ ancilla qubits, and $n$ input qubits such that:
  1. 1.
    if $x\in A_{\text{YES}}$ then $||\mathcal{V}^{(|x|)}_{x}||\geq c(|x|)$ ,
  2. 2.
    if $x\in A_{\text{NO}}$ then $||\mathcal{V}^{(|x|)}_{x}||\leq s(|x|)$ ,
where $\mathcal{V}^{(|x|)}_{x}$ is the acceptance operator of Eq. (2) corresponding to $V^{(|x|)}$ on input $x$.
Often in the literature the functions $c$ and $s$ are taken to be the constants $2/3$ and $1/3$, respectively. We choose this definition as it motivates the discussion of the promise gap $c(n)-s(n)$, which becomes important when defining #BQP in Section 2.3.
We will now recall the standard definition of BQP, and then afterwards define a complete problem for BQP that can be expressed in terms of properties of the acceptance operators.
Definition 2 (BQP).
Let $A=(A_{\text{YES}},A_{\text{NO}})$ be a promise problem and let $c,s:\mathbb{N}\to[0,1]$ be functions satisfying $c(n)-s(n)=\Omega(\text{poly}(n)^{-1})$. Then $A\in{\sf BQP}$ if and only if there exists a polynomial-time uniform family of quantum circuits $\{V^{(n)}:n\in\mathbb{N}\}$ acting on $a(n)=O(\text{poly}(n))$ ancilla qubits and $n$ input qubits such that:
  1. 1.
    if $x\in A_{\text{YES}}$ then $\bra{x,0^{a}}V^{(|x|)\dagger}\Pi_{\textup{out}}V^{(|x|)}\ket{x,0^{a}}\geq c(|x|)$ ,
  2. 2.
    if $x\in A_{\text{NO}}$ then $\bra{x,0^{a}}V^{(|x|)\dagger}\Pi_{\textup{out}}V^{(|x|)}\ket{x,0^{a}}\leq s(|x|)$
The BQP-complete problem we introduce related to verifier circuits is the decision problem of estimating their average accept probability.
Definition 3 (Average Accept Probability of Verifier Circuits).
Given functions $c,s:\mathbb{N}\to[0,1]$ satisfying $c(n)-s(n)=\Omega(\text{poly}(n)^{-1})$ and a polynomial-time uniform family of circuits $\{V^{(n)}:n\in\mathbb{N}\}$, each acting on $w(n)=O(\text{poly}(n))$ witness qubits, $a(n)=O(\text{poly}(n))$ ancilla qubits, and $n$ input qubits, the Average Accept Probability of Verifier Circuits problem is to, given an input $x$, decide whether
  1. 1.
    $\frac{\text{Tr}[\mathcal{V}^{(|x|)}_{x}]}{2^{w(|x|)}}\geq c(|x|)$ ,
  2. 2.
    or $\frac{\text{Tr}[\mathcal{V}^{(|x|)}_{x}]}{2^{w(|x|)}}\leq s(|x|)$ ,
promised one of these is the case, where $\mathcal{V}^{(|x|)}_{x}$ is the acceptance operator of Eq. (2) corresponding to $V^{(|x|)}$ on input $x$.
We show in Appendix B that this problem is BQP-complete.
[…]
$\mathcal{D}_{x}=(\bra{0^{a}}\otimes\mathbb{I}_{W})D_{x}^{\dagger}\Pi_{\text{out}}D_{x}(\ket{0^{a}}\otimes\mathbb{I}_{W})\,.$ (3)
With this we formally define DQC1.
Definition 4 (DQC1).
Let $A=(A_{\text{YES}},A_{\text{NO}})$ be a promise problem and let c,s: $\mathbb{N}\to[0,1]$ be functions satisfying $c(n)-s(n)=\Omega(\text{poly}(n)^{-1})$. Then $A\in{\sf DQC}_{1}$ if and only if for every input $x\in A$ there exists a polynomial-time generated quantum circuit $D_{x}$ with corresponding acceptance operator $\mathcal{D}_{x}$ as in Eq. (3) acting on $w(|x|)=O(\text{poly}(|x|))$ witness qubits and $a(|x|)=O(log(|x|))$ ancilla qubits such that:
  1. 1.
    If $x\in A_{\text{YES}}$ then $\frac{\textup{Tr}[\mathcal{D}_{x}]}{2^{w(|x|)}}\geq c(|x|)$ ,
  2. 2.
    If $x\in A_{\text{No}}$ then $\frac{\textup{Tr}[\mathcal{D}_{x}]}{2^{w(|x|)}}\leq s(|x|)$
In the literature DQC1 is often alternatively viewed as the class of problems that can be solved via sampling from circuits of the form given above. A canonical complete problem from this point of view is that of estimating the trace of $2^{n}\times 2^{n}$ unitary matrices up to an additive error normalized by $2^{n}$ (see Definition 6). Problems that are not decision problems are often referred to as DQC1-complete if they are polynomial-time reducible to this problem. We will adopt a similar point of view in this paper, although when showing DQC1-hardness of a problem we will only show that it holds for the decision problem version, noting that the use of a binary search can establish polynomial-time reducibility between the decision problem version of DQC1 and the problem of additively estimating traces of unitary matrices to the specified normalization.
2.3 #BQP
Some care needs to be taken in defining the counting class associated to QMA, known as #BQP. Intuitively, one would like to generalize the notion of computing the size of the set $R(x)$ (1) to the quantum case. This is not straightforward, as the number of possible quantum witnesses is uncountably infinite. Instead, a counting problem can be formulated as determining the dimension of the subspace of good witnesses. This subspace will be defined via the spectral properties of the acceptance operator $\mathcal{V}_{x}$. We first define the maps $N_{\geq a}$ from Hermitian operators to integers via
$N_{\geq a}(H)=|\{b:b\in\text{spec}(H),b\geq a\}|\,.$ (4)
That is, they return the number of eigenvalues of $H$ that are greater than or equal to $a$. One can then attempt to define #BQP via counting functions such that given a circuit $V$ and an input $x$ corresponding to an acceptance operator $\mathcal{V}_{x}$, return $N_{\geq c}(\mathcal{V}_{x})$. This ends up being too powerful, as the ability to return this would allow one to solve problems in the variant of QMA where the promise gap, $c-s$, is allowed to be exponentially small, which characterizes the class PSPACE [10]. Thus, one must be careful with how the promise gap is handled.
The original formulation of #${\sf\#BQP}$ [5, 21] circumvents this by imposing a promise that acceptance operators $\mathcal{V}_{x}$ corresponding to allowed problem instances $x$ do not have any eigenvalues in the interval $(c,s)$. Conveniently, this also guarantees that the output of any #${\sf\#BQP}$ function is unique but has the consequence that proving inclusion of problems in #${\sf\#BQP}$ requires adding extra promises. In particular, Ref. [5] showed that the density of states is a #${\sf\#BQP}$-complete problem but only for Hamiltonians that satisfy a potentially nonphysical promise on their spectrum. As Ref. [5] notes, there is an alternative approach to defining #${\sf\#BQP}$ where the output is allowed to “miscount” eigenvalues in the interval $(c,s)$. This was later generalized to define the notion of relative-error approximations to #${\sf\#BQP}$ problems and to show that this is equivalent to problems such as computing relative-error approximations to quantum partition functions under polynomial-time reductions [4]. Here, we choose to work with the miscounting definition of #${\sf\#BQP}$ because it does not require extra assumptions on the associated verifier circuits.
In order to rigorously define #BQP problems with the allowed miscounting in the promise gap we will define them as relations $r\subset J\times\mathbb{Z}^{+}$ between a set of inputs $J\subset\{0,1\}^{*}$, and the set of positive integers $\mathbb{Z}^{+}$. We use the lower case $r$ to distinguish them from the relations $R$ we used to define #P. Note that this definition includes standard functions as a special case of relations satisfying the property that every $x\in J$ has exactly one $y\in\mathbb{Z}^{+}$ related to it. We denote the set of outputs related to the input $x\in J$ via
$r(x)\equiv\{y:r(x,y)\}\,.$ (5)
A relational problem is then a problem where given an input $x\in J$, the goal is to return any $y\in r(x)$. With this in mind we can now define #BQP.
Definition 5 (#BQP).
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if there exists a uniformly generated family of quantum circuits $V^{r,n}$ and functions $c^{r},s^{r}:\mathbb{N}\to[0,1]$ satisfying $c^{r}(n)-s^{r}(n)=\Omega(\text{poly}(n)^{-1})$ such that
$r(x)=[N_{\geq c^{r}(|x|)}(\mathcal{V}_{x}^{r,|x|})..N_{\geq s^{r}(|x|)}(\mathcal{V}_{x}^{r,|x|})]\,.$
For the purposes of this work, certain parameters of the circuits $V^{r}_{x}$ will be very important. These parameters are defined as follows: given a family of verifier circuits $\{V^{r,n}\}_{n\in\mathbb{N}}$ in the $\{\text{Toffoli},H,S\}$ gate set where $r$ denotes the family and $n$ denotes the input size, define the parameters $w_{r},a_{r},t_{r},h_{r}:\mathbb{Z}^{+}\to\mathbb{Z}^{+}$ for a given input size $n$ via
  • $w_{r}(n)$ as the number of qubits in the witness register,
  • $a_{r}(n)$ as the number of qubits in the ancilla register,
  • $t_{r}(n)$ as the total number of gates,
  • $h_{r}(n)$ as the total number of $H$ gates.
2.4 Additive Error Approximations
In Ref. [1] it was shown that all #P and GapP functions have a natural level of additive approximability related to the length of the witnesses of the associated NP problems defining them. In particular, given a #P or GapP function $f$ with witnesses on $w(|x|)$ bits for inputs of size $|x|$, the authors show that there exists a probabilistic algorithm on input $x$ that outputs an estimate $\tilde{f}(x)$ satisfying
$|f(x)-\tilde{f}(x)|\leq\epsilon u(|x|)\,,$ (6)
to a normalization $u(|x|)=2^{w(|x|)}$ with probability $1-\delta$ that runs in time polynomial in $1/\epsilon$ and $\log(1/\delta)$. The fact that this level of additive approximation is equal complexity for both #P and GapP follows immediately from another result of the paper, namely that additive approximations with a normalization $u$ are closed under subtraction.
Here we will study whether the complexity of analogous additive approximations to #BQP relations are of a different complexity. In order to do so, we adapt the definitions of additive approximations from Ref. [1] to relational problems. We also distinguish between the problem of returning an additive approximation to a relation and algorithms for doing so.
Definition 6 (Additive Error Approx.).
Given a relation $r\subset J\times\mathbb{Z}^{+}$ and a normalization $u:\mathbb{Z}^{+}\to\mathbb{R}^{+}$, the problem of returning an additive error approximation to $r$ with normalization $u$ to precision $\epsilon>0$ for any input $x\in J$, is to return a number $\tilde{r}(x,u)$ satisfying
$\min\{|\tilde{r}(x,u)-y|:y\in r(x,\cdot)\}\leq\epsilon u(|x|)\,.$
Intuitively this just means that the estimate needs to be close to at least one of the output $y$’s related to an input $x$. Using this definition we can then formally define the additive approximation algorithms we will be considering throughout this work.
Definition 7 (Additive Approx. Algorithm).
A classical (quantum) additive approximation algorithm for a relation $r\subset J\times\mathbb{Z}^{+}$ with normalization $u$ is a probabilistic (quantum) algorithm that given an input $x\in J$ returns a number $\tilde{r}(x,u)$ satisfying the additive error approximation requirements of Definition 6 with probability at least $1-\delta$ and runs in time (and size) polynomial in $|x|$, $1/\epsilon$ and $\log(1/\delta)$.
For the quantum algorithms, by polynomial time and size we mean polynomially many runs of polynomial-size quantum circuits and polynomial-time classical pre- and post-processing is allowed.
2.5 The Complexity of Additive Error Approximations to #BQP
Here we present results related to additively approximating #BQP relations. The proofs for all of the theorems in this section are contained in Section 3.1. For our first result, we demonstrate that quantum additive approximation algorithms exist for all #BQP functions with normalization factors $u$ exponential in the number of qubits of the witness space.
Theorem 2.1.
For all relations $r\in$BQP with corresponding acceptance operators $\mathcal{V}_{x}^{r,|x|}$ on input $x$, there are quantum additive-error approximation algorithms with normalization factor $u(|x|)=2^{w_{r}(|x|)}$.
This result is analogous to the results in Ref. [1], but with the normalization now being determined by the number of qubits rather than the number of bits describing the witnesses. We next show that this level of additive approximation is in a sense the best possible.
Theorem 2.2.
The problem of returning additive-error approximations for all #BQP relations $r$ with corresponding acceptance operators $\mathcal{V}_{x}^{r,|x|}$ on input $x$ to normalization $u(|x|)=2^{cw_{r}(|x|)}$ for positive $c<1$ satisfying $1-c=\Omega(\text{poly}(|x|)^{-1})$ is #BQP-hard.
Our next result implies that it is unlikely that there exists classical additive approximation algorithms for #BQP functions with the same normalization $u$ of Theorem 2.1.
Theorem 2.3.
Returning additive-error approximations with $u(|x|)=2^{w_{r}(|x|)}$ for all #BQP relations $r$ with corresponding acceptance operators $\mathcal{V}_{x}^{r,|x|}$ on input $x$ is BQP-hard.
This suggests that this level of approximation can not be achieved by any polynomial-time classical algorithm, and together with Theorem 2.1 suggests that this level of additive approximation to #BQP relations characterizes the difficulty polynomial time quantum computation.
We next turn to looking for classical additive approximation algorithms to #BQP relations. Given the fact that exact counting classes are equivalent under polynomial-time reductions [5, 11], and that as shown in Ref. [1] there are efficient classical additive error approximations to all #P and GapP problems to some normalization $u$ there must exist classical additive approximations to #BQP relations. The key point that makes this consistent with the previous BQP-hardness result is that the reductions between the classes have a profound effect on the size of the witnesses, leading to a very different normalization $u$ between the quantum and classical approximation algorithms. We use the reduction in Ref. [5] that maps a #BQP relation to a GapP function with a witness size that results in a normalization $u$ which is significantly larger than the one in Theorem 2.1. A downside of the construction is that the achievable normalization likely depends on the gate set used.
Theorem 2.4.
For a #BQP relation $r$ with input $x$, let $h_{r}\equiv h_{r}(|x|)$ be number of Hadamard gates in the circuit $V^{r,|x|}$ using the $\{\text{Toffoli},H,S\}$ gate set and let $t_{r}\equiv t_{r}(|x|)$ be the depth and $a_{r}\equiv a_{r}(|x|)$ be the number of ancilla qubits. Then classical additive error approximations exist with normalization factor $u(|x|)=2^{2t_{r}(w_{r}+a_{r})-(a_{r}+1)-h_{r}}$.
Note that this is in general not a particularly useful approximation, as $u$ is typically much greater than the range of values that the #BQP function can return in the first place.
2.6 The Relationship between Additive Error Approximations to #BQP and DQC1
We now turn to analyzing the connection between additive approximations to #BQP relations and the complexity class DQC1. The proofs of the theorems in this section are presented in Section 3.2. This connection is motivated by a collection of results showing that additive approximations to certain problems related to or contained in #BQP, including the log-local quantum partition function and the log-local quantum density of states problems, are DQC1-complete [2, 9, 15]. Although the density of states problem is #BQP-complete, the reductions used affect the levels of additive error achievable, so conclusions are not immediate about the relationship between additive approximations to general #BQP relations and DQC1.
We now elucidate the connection between #BQP generally and DQC1 by showing that additive approximations to a subset of #BQP relations, namely those which have verifier circuits requiring only logarithmically many ancilla qubits, have the same computational power as DQC1. In order to show this we first need to define a version of #BQP with the notion of uniform generation that is used in DQC1.
Definition 8 (#BQP Alternative Version).
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if for each input $x\in J$ there exists a polynomial-time generated quantum circuit $D_{x}^{r}$ and functions $c^{r}$,$s^{r}:\mathbb{N}\to[0,1]$ satisfying $c^{r}(n)-s^{r}(n)=\Omega(\text{poly}(n)^{-1})$ such that
$r(x)=[N_{\geq c^{r}(|x|)}(\mathcal{D}_{x}^{r})..N_{\geq s^{r}(|x|)}(\mathcal{D}_{x}^{r})].$
Our first result demonstrates that one only needs additive approximations to the subclass of #BQP relations with ancillas that scale logarithmically in the size of the number of witness qubits in order to solve DQC1 problems.
Theorem 2.5.
Returning additive-error approximations with $u(|x|)=2^{w_{r}(|x|)}$ for #BQP relations $r$ with corresponding verifier circuits $D_{x}^{r}$ with $a_{r}(|x|)=O(\log(w_{r}(|x|)))$ is DQC1-hard.
We next show that the ability to solve DQC1 decision problems a polynomial number of times allows for the computation of additive-error approximations to #BQP relations with logarithmically many ancillas.
Theorem 2.6.
The ability to solve DQC1 problems can be used to provide additive-error approximations with $u(|x|)=2^{w_{r}(|x|)}$ to all #BQP relations $r$ with corresponding verifier circuits $D_{x}^{r}$ on input $x$ such that $a_{r}(|x|)=O(\log(w_{r}(|x|)))$.
An immediate corollary of Theorems 2.5 and 2.6 is that returning additive approximations to #BQP relations with logarithmically many ancillas is complete for the alternative non-decision formulation of DQC1 described in Section 2.2.
3 Proofs
3.1 Additive Error Approximations to #BQP
In this section we provide proofs for Theorems 2.1, 2.2, 2.3 and 2.4.
Quantum Additive Approximation Algorithms to #BQP relations.
Here we prove Theorem 2.1.
Proof.
Given a verifier circuit $V^{r}_{x}$ corresponding to input $x$ acting on $w_{r}(|x|)$ witness qubits and $a_{r}(|x|)$ ancilla qubits with completeness and soundness parameters $c$ and $s$, we construct a new circuit $\tilde{V}^{r}_{x}$ with amplified completeness and soundness parameters which will be chosen later. Note that such a circuit can always be constructed while preserving the dimension of the witness subspace [19, 5]. The algorithm is then to generate $M=\text{poly}(|x|)$ uniformly random computational basis states for the witness subspace and send them through the new verifier circuit, and then count how many accept, and multiply this number by $2^{w_{r}(|x|)}/M$. Let $X_{i}$ be the indicator function that evaluates to $1$ if the $i$-th bit string accepts. Then this estimator is
$X=\frac{2^{w_{r}(|x|)}}{M}\sum_{i=1}^{M}X_{i}\,.$ (7)
It can be shown that $\mathbb{E}[X]=\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]$ for the amplified version of the acceptance operator in Eq. (2), which due to the new $\tilde{c}$ and $\tilde{s}$ parameters satisfies
$N_{\geq c}(\mathcal{V})-(1-\tilde{c})2^{w_{r}(|x|)}\leq\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]\leq N_{\geq s}(\mathcal{V})+\tilde{s}2^{w_{r}(|x|)}\,.$ (8)
The variance of $X$ is then given by
$\text{Var}[X]=\frac{1}{M}\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}](2^{w_{r}(|x|)}-\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}])\,.$ (9)
From this it follows that
$\textbf{Pr}(|X-\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]|\geq\epsilon 2^{w_{r}(|x|)})\leq$ $\frac{\text{Var}[X]}{\epsilon^{2}2^{2w_{r}(|x|)}}$ (10)
$=$ $\frac{1}{M\epsilon^{2}}\frac{\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]}{2^{w_{r}(|x|)}}\left(1-\frac{\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]}{2^{w_{r}(|x|)}}\right)$ (11)
$\leq$ $\frac{1}{M\epsilon^{2}}\,.$ (12)
Thus with constant failure probability $\delta\leq 1/4$ and $M=O(1/\epsilon^{2})$ samples we can output a sample $\tilde{X}$ satisfying
$N_{\geq c}(\mathcal{V})-(1-\tilde{c}+\epsilon)2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s}(\mathcal{V})+(\epsilon+\tilde{s})2^{w_{r}(|x|)}\,.$ (13)
Now, we choose $\tilde{c}\geq 1-1/\text{poly}(|x|)$ and $\tilde{s}\leq 1/\text{poly}(|x|)$ such that both $(1-\tilde{c}+\epsilon)$ and $\tilde{s}+\epsilon$ are less than $2\epsilon=\epsilon^{\prime}=\Omega(\text{poly}(|x|)^{-1})$ so that the sample satisfies
$N_{\geq c}(\mathcal{V})-\epsilon^{\prime}2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s}(\mathcal{V})+\epsilon^{\prime}2^{w_{r}(|x|)}\,.$ (14)
The success probability can then be amplified to $1-\delta^{\prime}$ with $\log(1/\delta^{\prime})$ applications of the previous procedure and then taking the median of the $\log(1/\delta^{\prime})$ estimates. We thus have a quantum algorithm that outputs an additive error estimate with normalization $u(|x|)=2^{w_{r}(|x|)}$ that has a total runtime and size polynomial in $|x|$, $1/\epsilon^{\prime}$ and $\log\left(1/\delta^{\prime}\right)$. ∎
Better approximation algorithms are #BQP-hard.
Next we prove Theorem 2.2 adapting a proof from Ref. [23].
Proof.
Given an input $x$ to a #BQP relation $r$ with corresponding verifier circuit $V_{x}^{r}$ and completeness and soundness parameters $c^{r}(|x|)$ and $s^{r}(|x|)$, define a new verifier circuit $V_{x}^{r^{\prime}}=V_{x}^{r}\otimes\mathbb{I}^{\otimes l}$ corresponding to another #BQP relation $r^{\prime}$ where the witness subspace has been expanded by $l$ qubits where $l$ will be chosen later. It then follows that $N_{\geq c^{r}(|x|)}(\mathcal{V}^{{}^{\prime}r}_{x})=2^{l}N_{\geq c^{r}(|x|)}(\mathcal{V}^{r}_{x})$ and $N_{\geq s^{r}(|x|)}(\mathcal{V}^{{}^{\prime}r}_{x})=2^{l}N_{\geq s^{r}(|x|)}(\mathcal{V}^{r}_{x})$. Now, assume an additive error estimate $\tilde{r}^{\prime}(x,u)$ to the #BQP instance corresponding to the circuit $V_{x}^{r^{\prime}}$ can be returned with a normalization $u_{r^{\prime}}(|x|)=2^{c(w_{r}(|x|)+l)}$ to some inverse polynomial precision $\epsilon<1$. Then the estimate $\tilde{r}=\tilde{r}^{\prime}(x,u)/2^{l}$ satisfies
$\min\{|\tilde{r}-y|:y\in r(x,\cdot)\}\leq 2^{w_{r}(|x|)-(1-c)l}\,.$ (15)
If we take $l>\frac{w_{r}(|x|)}{1-c}+1$ then the error is less than $1/2$ and by rounding to the nearest integer, a number in $r(x,\cdot)$ can always be returned. This takes $l=O\left(\frac{w_{r}(|x|)}{1-c}\right)$ extra qubits and thus $l$, and the size of $V^{r^{\prime}}_{x}$, will be upper bounded by a polynomial in $|x|$ whenever $1-c=\Omega(\text{poly}(|x|)^{-1})$.
Additive approximations to #BQP are BQP-hard
Prior to proving Theorem 2.3, we introduce a different version of #BQP that allows for the $c^{r}$ and $s^{r}$ functions to depend on the input $x$, rather than just its size, as utilizing this version will greatly simplify the proof.
Definition 9 (#BQP${}_{(c^{r},s^{r})}$).
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if there exists a uniformly generated family of quantum circuit $V^{r,n}$ and a polynomial $q^{r}$ such that
$r(x)=[N_{\geq c^{r}(x)}(\mathcal{V}_{x}^{r,|x|})..N_{\geq s^{r}(x)}(\mathcal{V}_{x}^{r,|x|})]\,,$
where $c^{r},s^{r}:\{0,1\}^{*}\to[0,1]$ are efficiently computable functions satisfying $c^{r}(x)-s^{r}(x)>1/q^{r}(|x|)$.
We show in Appendix A that Definition 9 is equivalent to Definition 5 in order to ensure that this new definition is not more powerful. This yields the following lemma.
Lemma 3.1.
#BQP${}_{(c^{r},s^{r})}$ = #BQP
We are now in a position to prove Theorem 2.3.
Proof of Theorem 2.3.
The goal here is to show that if we can solve a #BQP${}_{(c^{r},s^{r})}$ relation $r$ on a series of polynomially many inputs $x_{i}$ to additive error, with normalizations $u_{i}=2^{w(|x_{i}|)}$, then we can solve an input $x$ of an instance of the Average Accept Probability of Verifier Circuits problem in Definition 3 with parameters $c(|x|)$ and $s(|x|)$. In particular, let $x\mapsto x_{i}=y_{i}x$ where $y_{i}$ is a bit string which encodes the correct completeness and soundness parameters to use for the corresponding input. Choosing a fixed length prefix-free encoding for $y_{i}$ ensures uniqueness of each concatenated bit string, and each $x_{i}$ has the same number of bits, which will be necessary to avoid repeated inputs. Each input $x_{i}$ to the #BQP${}_{(c^{r},s^{r})}$ relation we are constructing will have a verifier circuit $V^{|x_{i}|}$ whose corresponding acceptance operator will be $\mathcal{V}_{x}$, the same operator as that of the Average Accept Probability of Verifier Circuits input. The $y_{i}$ part of the bit string will then determine the $c_{i}$ and $s_{i}$ interval.
Then we begin by partitioning the interval $[0,1]$ into $M=\text{poly}(|x|)$ subintervals $[s_{i},c_{i}]$, each of size $1/\text{poly}(|x|)$, corresponding to the different completeness and soundness parameters specified by $y_{i}$. Specifically, we consider the subintervals $[s_{i},c_{i}]$ where $c_{i}=(M-i)/M$ and $s_{i}=c_{i}-1/4M$ for $11\leq i\leq M-$ 1. For $c_{M}$ and $s_{M}$ the oracle will always return $2^{w(|x|)}$. Then querying the #BQP${}_{(c^{r},s^{r})}$ oracle for each problem input $x_{i}$ returns an estimate
$\hat{n}_{i}=N_{\geq c_{i}}+\delta_{i}+\varepsilon_{i}\,,$ (16)
with $|\delta_{i}|\in[0,N_{[s_{i},c_{i}]}]$ and $|\varepsilon_{i}|\in[0,\epsilon 2^{w(|x_{i}|)}]$. Letting $\hat{n}_{0}=0$, we define the two quantities $\hat{\chi}_{i}=\hat{n}_{i}-\hat{n}_{i-1}$ and $\Delta_{i}=c_{i}+1/2M$. Then we have
$\sum_{i=1}^{M}\Delta_{i}\hat{\chi}_{i}$ $=\sum_{i=1}^{M}\Delta_{i}\left(N_{[c_{i},c_{i-1}]}+\delta_{i}-\delta_{i-1}+\varepsilon_{i}-\varepsilon_{i-1}\right)\,,$ (17)
$=\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}+\frac{1}{M}(\delta_{i}+\varepsilon_{i})\,.$ (18)
We will now show that this quantity approximates the trace of the associated acceptance operator $\mathcal{V}_{x}$ of the BQP problem. First, note that
$|\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}_{[c_{i},c_{i-1}]}(\mathcal{V}_{x^{\prime}})|\leq\frac{1}{2M}N_{[c_{i},c_{i-1}]}\,,$ (19)
which then implies that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}(\mathcal{V}_{x})\bigg{|}\leq\frac{2^{w(|x|)}}{2M}\,,$ (20)
where we have used the fact that $2^{w(|x_{i}|)}=2^{w(|x|)}$ for all $i$, or rather, the size of the witness subspace for each input $x_{i}$ is the same size as that of $x$ even though the classical inputs may vary in size.
Finally, we bound the error resulting from $\delta_{i}$ and $\varepsilon_{i}$ by noting that
$\frac{1}{M}\sum_{i=1}^{M}\delta_{i}\leq\frac{2^{w(|x|)}}{M},\quad\text{and}\quad\sum_{i=1}^{M}\varepsilon_{i}\leq\epsilon 2^{w(|x|)}\,,$ (21)
where the first inequality holds from the fact that $[s_{i},c_{i}]\cap[s_{j},c_{j}]=\emptyset$ for all $i\neq j$. Setting $\epsilon=1/M$, we add the errors from each contribution and find that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}\hat{\chi}_{i}-\text{Tr}(\mathcal{V})\bigg{|}\leq\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{2M}=\frac{5}{2}\frac{2^{w(|x|)}}{M}\,.$ (22)
Choosing $M>5/4(c(|x|)-s(|x|))$ suffices to distinguish ebetween the accepting and rejecting cases of Definition 3, a BQP-complete problem. Thus, returning additive approximations to #BQP problems is BQP-hard.
Classical additive approximations to #BQP relations
Here we prove Theorem 2.4.
Proof of Theorem 2.4.
Given a #BQP relation $r$ and an input $x$ with a corresponding verifier circuit $V_{x}^{r}$, set $n=w_{r}(|x|)$ as the number of witness qubits, $a=a_{r}(|x|)$ as the number of ancilla qubits, $T=t_{r}(|x|)$ as the size of the circuit in the $\{H,S,\text{Toffoli}\}$ gate set, and $h=h_{r}(|x|)$ as the number of Hadamard gates. Let $V_{x}^{r}=g_{T}\cdots g_{1}$ in the chosen gate set and let $\mathcal{V}$ be the corresponding verifier operator. First note that because of the completeness and soundness parameters we have that
$N(\mathcal{V})_{\geq c}-(1-c)2^{n}\leq\text{Tr}[\mathcal{V}]\leq N(\mathcal{V})_{\geq s}+s2^{n}\,.$ (23)
Define the operators $q_{i}$ via $q_{i}=g_{i}$ if $g_{i}\in\{\text{Toffoli},S\}$ and $q_{i}=\sqrt{2}g_{i}$ if $g_{i}=H$ so that all of the matrix elements of $q_{i}$ are in $\{1,-1,i,-i\}$. Now expand the trace
[…]
$=$ $\sum_{x,y,z_{j}}\langle 0^{a}x|g_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|g_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|g_{T}^{\dagger}|1y\rangle\!\langle 1y|g_{T}|z_{T}\rangle\!\langle z_{T}|g_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|g_{1}|0^{a}x\rangle$ (25)
$=$ $\frac{1}{2^{h}}\sum_{x,y,z_{j}}\langle 0^{a}x|q_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|q_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|q_{T}^{\dagger}|1y\rangle\!\langle 1y|q_{T}|z_{T}\rangle\!\langle z_{T}|q_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|q_{1}|0^{a}x\rangle.$ (26)
We can then define two #P functions $g$ and $f$ as follows: for $g$ count the number of paths $p\in\{(x,y,z_{1},...,z_{2(T+1)})\}$ such that the product
$\langle 0^{a}x|q_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|q_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|q_{T}^{\dagger}|1y\rangle\!\langle 1y|q_{T}|z_{T}\rangle\!\langle z_{T}|q_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|q_{1}|0^{a}x\rangle$ (27)
evaluates to $1$ and for $f$ count the number of paths such that it evaluates to $-1$. It then follows that the GapP function defined by $r=g-f$ satisfies
$\frac{r}{2^{h}}=\text{Tr}[\mathcal{V}]\,.$ (28)
Next, it follows from Ref. [1] that we can additively approximate $r$ with $u=2^{N^{*}}$ where $N^{*}=|p|$ is the number of bits needed to describe the paths $p$. Since $|x|=n$, $|y|=n+a-1$, and $|z_{j}|=n+a$ for all $j$, it follows that the number of bits to describe the path is $N^{*}=2T(n+a)-(a+1)$. Now, it follows by approximating $r$ and dividing by $2^{h}$, we get an additive approximation $\tilde{\mathcal{V}}$ to $\text{Tr}[\mathcal{V}]$ with $u=2^{N^{*}}/2^{h}$ we have
$N_{\geq c}(\mathcal{V})-(1-c)2^{n}-\epsilon 2^{N^{*}-h}\leq\text{Tr}[\tilde{\mathcal{V}}]\leq N_{\geq s}(\mathcal{V})+s2^{n}+\epsilon 2^{N^{*}-h}$ (29)
which is equivalent to
$N_{\geq c}(\mathcal{V})-\epsilon^{\prime}2^{N^{*}-h}\leq\text{Tr}[\tilde{\mathcal{V}}]\leq N_{\geq s}(\mathcal{V})+\epsilon^{\prime\prime}2^{N^{*}-h}$ (30)
when
$\epsilon^{\prime}=(1-c)2^{n-N^{*}+h}+\epsilon\,,\qquad\epsilon^{\prime\prime}=s2^{n-N^{*}+h}+\epsilon\,.$ (31)
Since $n-N^{*}+h=n+a+h+1-2T(n+a)$ and $h\leq T$, $\epsilon^{\prime}$ and $\epsilon^{\prime\prime}$ are always inverse polynomial whenever $\epsilon$ is. Thus there are efficient classical additive approximations for #BQP functions when $u=2^{2T(n+a)-(a+1)-h}$. ∎
3.2 #BQP and DQC1
In this section we prove Theorems 2.5 and 2.6, noting that the proof of Theorem 2.5 follows closely to that of Theorem 2.3.
Additive approximations to #BQP with logarithmic ancillas is DQC1-complete.
Proof of Theorem 2.5.
Again, just as in the proof of Theorem 2.3, the goal here is to show that if we can solve a #BQP${}_{(c^{r},s^{r})}$ relation $r$ on a polynomial number of inputs $x_{i}$ to additive error, with normalizations $u_{i}=2^{w(|x_{i}|)}$, then we can solve the Average Accept Probability of Verifier Circuits problem in Definition 3. However, there are a few notable differences here. First, the notion of uniform generation of quantum circuits changes when considering DQC1 so rather than starting from Definition 5, we use the alternative definition of #BQP in Definition 8. Its equivalence to #BQP${}_{(c^{r},s^{r})}$ follows exactly that of the proof in Appendix A. Additionally, here we consider only a sublcass of #BQP problems, namely those whose verifier circuits $D^{|x|}$ require at most a logarithmic number of ancilla qubits. Provided each step in our reduction to the Average Accept Probability of Verifier Circuits problem, which is DQC1-complete in this regime, does not use more than $O(\log(|x|)$ ancilla qubits, then we will show that additively approximating any relation $r$ in this subclass of #BQP is DQC1-hard.
The techniques used in the proof Theorem 2.3 will be the exact same here, so rather than restating the proof, we will just ensure each step preserves the size of the ancilla space. In fact, the only part of the proof that can change the ancilla space is in the reduction of Definition 8 to Definition 9 in Appendix A, in which we employ the quantum singular value transformation. However, this step only requires a single ancilla qubit to implement. Thus, provided that the verifier circuits $D^{|x|}$ require only $O(\log(|x|)$ ancilla qubits, then a polynomial number of calls to the #BQP${}_{(c^{r},s^{r})}$ oracle to additive error allows us to additively approximate any DQC1 problem. ∎
Proof of Theorem 2.6.
Let $V_{x}^{r}$ be a given verifier circuit for a #BQP relation $r$ on an input $x$ with $w_{r}(|x|)$ qubits and $a_{r}(|x|)=O(\log(w_{r}(|x|)))$ ancilla qubits with completeness and soundness parameters $c(|x|)$ and $s(|x|)$ and an associated verifier operator $\mathcal{V}_{x}$. Then one can always construct a block encoding $A\coloneqq\bra{1}V_{x}^{r}(\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$, and then make use of Ref. [13, Theorem 19] to implement a singular value threshold projector on $A$. This allows us to construct a new block encoding $A_{\Phi}=\bra{1}V_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$, for some $\Phi\in\mathbb{R}^{p}$, and where $V_{\Phi}$ is given by the circuit in Fig. 1. The new encoding $A_{\Phi}$ now has singular values $\tilde{\varsigma}_{i}$ which are $\varepsilon$-close to 00 or $1$ depending on whether the singular values $\varsigma_{i}$ of $A$ are below or above a threshold $t\in(0,1)$.
To construct the new block encoding $A_{\Phi}$, define a function on the interval $[-1,1]$ which takes the value 00 on the interval $[-t,t]$ and $1$ elsewhere. A polynomial approximation $P$ can be constructed such that
  1. 1.
    $|P(x)|\leq 1$ for $x\in[-1,1]$ ,
  2. 2.
    $P(x)\in[1-\varepsilon,1]$ for $x\in[-1,-t-\Delta]\cup[t+\Delta,1]$ with $\Delta\in(0,1)$ ,
  3. 3.
    $P(x)\in[0,\varepsilon]$ for $x\in[-t+\Delta,t-\Delta]$  .
Additionally, it can be shown that this polynomial approximation is even and has degree $p=O(\log(1/\varepsilon)/\Delta)$ using techniques similar to that of Ref. [13, Corollary 16]. Finally, according to Ref. [13, Corollary 11], given the polynomial approximation $P$, one can construct a vector $\Phi\in\mathbb{R}^{p}$ which applies $P$ to the singular values of $A$, resulting in the new encoding $A_{\Phi}=P(A)$. The implementation of $P(A)$ is based on quantum signal processing techniques, and requires only one ancilla qubit with $p$ applications of $V_{x}^{r}$ and $(V_{x}^{r})^{\dagger}$ and $O(p)$ other elementary gates, as shown in Fig. 1.
Thus, given the verifier circuit $V_{x}^{r}$, we define the new verifier circuit $V_{\Phi}$ in Fig. 1 that implements the singular value threshold projector protocol on its block encoding $A$, setting $t=(s(|x|)+c(|x|))/2$ and $\Delta=(c(|x|)-s(|x|))/2$, to obtain the new encoding $A_{\Phi}=\bra{1}V_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$ with associated acceptance operator
$\mathcal{V}_{\Phi}$ $=A_{\Phi}^{\dagger}A_{\Phi}\,,$ (32)
[…]
$=\sum_{i}\tilde{\varsigma}_{i}\ket{\psi_{i}}\bra{\psi_{i}}\,,$ (34)
where |$\ket{\psi_{i}}$ form an orthonormal basis and $\tilde{\varsigma}_{i}$ are the amplified singular values such that $\tilde{\varsigma}_{i}\in[1-\varepsilon,1]$ when $\varsigma_{i}\geq c(|x|)$ and $\tilde{\varsigma}_{i}\in[0,\varepsilon]$ when $\varsigma_{i}\leq s(|x|)$. Due to the amplification, the trace of this operator satisfies
$N_{\geq c(|x|)}(\mathcal{V}_{x})-(2\varepsilon-\varepsilon^{2})2^{w_{r}(|x|)}\leq\text{Tr}[\mathcal{V}_{\Phi}]\leq N_{\geq s(|x|)}(\mathcal{V}_{x})+\varepsilon^{2}2^{w_{r}(|x|)}\,.$ (35)
Finally, the trace of this amplified acceptance operator can be additively estimated to normalization $u(|x|)=2^{w_{r}(|x|)}$ and precision $\epsilon$ via the same algorithm as in the proof of Theorem 2.1 using $O(1/\epsilon^{2})$ repetitions of $V_{\Phi}$. This outputs an estimate $\tilde{X}$ satisfying
$N_{\geq c(|x|)}(\mathcal{V}_{x})-(2\varepsilon-\varepsilon^{2}+\epsilon)2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s(|x|)}(\mathcal{V}_{x})+(\varepsilon^{2}+\epsilon)2^{w_{r}(|x|)}\,,$ (36)
with probability greater than $3/4$. Thus by taking $\varepsilon=O(\log(\text{poly}(\epsilon^{\prime})^{-1}))$ such that both $2\varepsilon-\varepsilon^{2}$ and $\varepsilon^{2}$ are less than or equal to $\epsilon$, an additive error estimate to normalization $2^{w_{r}(|x|)}$ and precision $\epsilon^{\prime}=2\epsilon$ can be obtained. Then, noting that $\Delta=(c(|x|)-s(|x|))/2=\Omega(\text{poly}(|x|)^{-1})$, an additive error estimate to normalization $u(|x|)=2^{w(|x|)}$ and to precision $\epsilon^{\prime}$ can be obtained using $O(\text{poly}\left(1/\epsilon^{\prime}\right))$ runs of $O(\log(\text{poly}(1/\epsilon^{\prime}))\cdot\text{poly}(|x|))$ size quantum circuits. Then repeating this procedure $\log(1/\delta^{\prime})$ times and taking the median of estimates can improve the success probability to $1-\delta^{\prime}$. Importantly since the circuits used in this algorithm require only a logarithmic overhead of $a_{r}(|x|)+2=O(\log(w_{r}(|x|)))$ pure ancilla qubits, they can be computed using DQC1 circuits.
Figure 1: The circuit $V_{\Phi}$ implementing the quantum singular value transformation $A_{\Phi}=P(A)$ given in Ref. [13] using a degree-$p$, even polynomial approximation. The block encodings $V_{x}^{r}$ and $V_{x}^{r\dagger}$ are applied alternately followed by applications of the reflection operators. In particular, controlled on the witness register being in the image of $\Pi$, the ancilla qubit is flipped, then a single qubit Pauli $Z$ rotation is applied to angles $\phi_{i}$, where the particular angles are classically determined according to the function being approximated by the polynomial. The final control gate performs an $X$ operation on the top qubit if and only if all of its control qubits are in the 00 state. The top $a_{r}(|x|)+2$ qubits are ancillas and the bottom $w_{r}(|x|)$ qubits are the witness qubits.
4 Discussion
The relationship to DQC1
Recall that the quantum density of states problem is #BQP-complete [5], and returning additive approximations to it to a specific $u$ is DQC1-complete [2, 8]. Therefore it is natural to wonder how this fact connects to our results. Here we elucidate on how the discrepancy in complexity between additive approximations to the density of states and additive approximations to #BQP functions arises. The key factor behind this discrepancy is how the size of witness space is affected by the reduction.
The reductions that show that the density of states is hard for #BQP (c.f. [5]) work by taking an $x$ input to a #BQP relation $r$ with an associated verifier circuit $V_{x}^{r}$ (note here we are using the alternative version of #BQP in Definition 8), and use the Feynman-Kitaev circuit to Hamiltonian mapping to construct a Hamiltonian whose low energy subspace corresponds to the subspace of witnesses that accept with high probability.